# MIMO Detection with Spatial Sigma-Delta ADCs: A Variational Bayesian Approach

### v1

@@ VI Numerical Results
Fig. 12 shows the SER of the LMMSE, $1^{\rm st}$-order SD-VB, and $2^{\rm nd}$-order SD-VB algorithms for $3$-bit $\Sigma\Delta$ architectures implemented with different steering angles $\phi=2\pi\frac{d}{\lambda}\sin(\theta)$. The steering angles are varied while the center angle of the sector is set as $\theta_{0}=0^{\circ},30^{\circ}$, and $60^{\circ}$ in Figs. 12(a), 12(b), and 12(c), respectively. Results for MF-VB are included as a reference, although its performance does not depend on $\phi$ since it does not use the $\Sigma\Delta$ implementation. As expected, the lowest SER for the $1^{\rm st}$- and $2^{\rm nd}$-order SD-VB architectures is obtained when the steering angle is chosen to match the center angle of the sector, i.e., $\theta=\theta_{0}$. The SER performance of all algorithms degrades as the center angle of the sector increases, since the ability of the array to spatial separate the signals decreases. LMMSE performs best without a phase offset, and degrades significantly as the angular sector of the users moves away from broadside.
VII Conclusions
@@
In this paper, we have developed two MIMO detection algorithms based on the VB inference framework in massive MIMO systems with few-bit $\Sigma\Delta$ ADCs. We first modeled the Bayesian networks for the $1^{\rm st}$- and $2^{\rm nd}$-order $\Sigma\Delta$ receivers, based on which the variational distributions of the latent variables were obtained in closed form. We then proposed two iterative algorithms for MIMO detection with efficient updates and low-complexity implementation. Simulation results showed that the proposed $1^{\rm st}$- and $2^{\rm nd}$-order SD-VB algorithms achieve their best data detection performance when the users are confined to angular sectors less than $80^{\circ}$ wide and with antenna spacings on the order of $1/4$ to $1/6$ a wavelength. Under these conditions, a $2^{\rm nd}$-order SD-VB approach with $3$-bit ADCs can outperform the MF-VB algorithm implemented with infinite resolution quantization, since SD-VB directly exploits the users’ spatial correlation. The simulation results also suggest using $1^{\rm st}$-order SD-VB rather than the $2^{\rm nd}$-order SD-VB implementation with $1$-bit quantization due to potential ADC overloading and unstable performance.
Appendix A Calculation of $\mathsf{F}_{r}(\mu,\gamma,a,b)$ and ${\sf G}_{r}(\mu,\gamma,a,b)$
@@
where $f(x)$ and $F(x)$ denote the PDF and CDF of a standard Gaussian random variable, respectively.
@@
Appendix B Calculation of ${\sf F}_{s}(\mathbf{z}_{k},\mathbf{h}_{k},\gamma)$ and ${\sf G}_{s}(\mathbf{z}_{k},\mathbf{h}_{k},\gamma)$
Given $\mathbf{z}_{k}=\mathbf{h}_{k}s_{k}+\mathcal{CN}(\mathbf{0},\gamma^{-1}\mathbf{I}_{N})$, the posterior distribution of $s_{k}$ given $\mathbf{z}_{k}$ is

### v2

@@ V Numerical Results
Fig. 15 shows the SER of the LMMSE, $1^{\rm st}$-order SD-VB, and $2^{\rm nd}$-order SD-VB algorithms for $3$-bit $\Sigma\Delta$ architectures implemented with different steering angles $\phi=2\pi\frac{d}{\lambda}\sin(\theta)$. The steering angles are varied while the center angle of the sector is set as $\theta_{0}=0^{\circ},30^{\circ}$, and $60^{\circ}$ in Figs. 15(a), 15(b), and 15(c), respectively. Results for MF-VB are included as a reference, although its performance does not depend on $\phi$ since it does not use the $\Sigma\Delta$ implementation. As expected, the lowest SER for the $1^{\rm st}$- and $2^{\rm nd}$-order SD-VB architectures is obtained when the steering angle is chosen to match the center angle of the sector, i.e., $\theta=\theta_{0}$. The SER performance of all algorithms degrades as the center angle of the sector increases, since the ability of the array to spatial separate the signals decreases. LMMSE performs best without a phase offset, and degrades significantly as the angular sector of the users moves away from broadside.
VI Conclusions
@@
In this paper, we have developed two MIMO detection algorithms based on the VB inference framework in massive MIMO systems with few-bit $\Sigma\Delta$ ADCs. We first modeled the Bayesian networks for the $1^{\rm st}$- and $2^{\rm nd}$-order $\Sigma\Delta$ receivers, based on which the variational distributions of the latent variables were obtained in closed form. We then proposed two iterative algorithms for MIMO detection with efficient updates and low-complexity implementation. Simulation results showed that the proposed $1^{\rm st}$- and $2^{\rm nd}$-order SD-VB algorithms achieve their best data detection performance when the users are confined to angular sectors less than $80^{\circ}$ wide and with antenna spacings on the order of $1/4$ to $1/6$ a wavelength. Under these conditions, a $2^{\rm nd}$-order SD-VB approach with $3$-bit ADCs can outperform the MF-VB algorithm implemented with infinite resolution quantization, since SD-VB directly exploits the users’ spatial correlation. The $2^{\rm nd}$-order SD-VB maintained superior SER performance under MC due to the fact it exploits correlation between signals at adjacent antennas for noise shaping and hence partially offsets the negative effects of channel correlation. The simulation results also suggest there are benefits to using $1^{\rm st}$-order SD-VB rather than the $2^{\rm nd}$-order SD-VB implementation with $1$-bit quantization due to potential ADC overloading and unstable performance. In future work, we will explore joint grid-less channel estimation and data detection using VB-based techniques for massive MIMO systems with few-bit $\Sigma\Delta$ ADCs.
Appendix A Linear Model for the $1^{\rm st}$-Order $\Sigma\Delta$ Architecture
From (12), the input to the quantizers of the $1^{\rm st}$-order $\Sigma\Delta$ architecture is expressed as
$\mathbf{r}=\mathbf{Ux}-\mathbf{Vy},$ (58)
where $\mathbf{V}=\mathbf{U}-\mathbf{I}_{N}$ and the matrix $\mathbf{U}$ is given by
We model the $1^{\rm st}$-order $\Sigma\Delta$ quantization operation as
$\mathbf{y}=\mathcal{Q}(\mathbf{r})=\bm{\Gamma}\mathbf{r}+\mathbf{q},$ (60)
where $\bm{\Gamma}$ is an $N\times N$ matrix, and $\mathbf{q}=\mathbf{y}-\bm{\Gamma}\mathbf{r}$ represents the effective quantization noise corresponding to the choice of $\bm{\Gamma}$. Following the approach in [18, 20], we design $\bm{\Gamma}$ using the Bussgang decomposition such that $\mathbb{E}[r_{i}q_{i}^{*}]=0$ for the corresponding pairs of elements in $\mathbf{r}$ and $\mathbf{q}$. This results in a diagonal $\bm{\Gamma}$ whose $i$-th diagonal element $\xi_{i}$ given by $\xi_{i}={\mathbb{E}[r_{i}y_{i}^{*}]}/{\mathbb{E}[|r_{i}|^{2}]}$. Substituting (58) into (60), we obtain
$\mathbf{y}$ $=\mathbf{\Gamma}\mathbf{Ux}-\mathbf{\Gamma}\mathbf{Vy}+\mathbf{q},$
$\Leftrightarrow\mathbf{y}$ $=(\mathbf{I}+\bm{\Gamma}\mathbf{V})^{-1}\bm{\Gamma}\mathbf{Ux}+(\mathbf{I}+\bm{\Gamma}\mathbf{V})^{-1}\mathbf{q}.$ (61)
Based on [20, 18], the quantization step size can be chosen such that $\xi_{i}=1$, $\forall i$, implying $\bm{\Gamma}=\mathbf{I}$. Using the identity $\mathbf{U}=\mathbf{V}+\mathbf{I}$, (A) can be simplified as
$\mathbf{y}=(\mathbf{I}+\mathbf{V})^{-1}\mathbf{Ux}+(\mathbf{I}+\mathbf{V})^{-1}\mathbf{q}=\mathbf{x}+\mathbf{U}^{-1}\mathbf{q}.$ (62)
Appendix B Calculation of $\mathsf{F}_{r}(\mu,\gamma,a,b)$ and ${\sf G}_{r}(\mu,\gamma,a,b)$
@@
where $F(x)$ and $f(x)$ are respectively the cumulative distribution function and probability density function of a logistic random variable $x$ with zero mean and unit variance, defined as $F(x)=1/(1+\mathrm{e}^{-3x/\sqrt{\pi}})$ and $p(x)=\frac{3}{\sqrt{\pi}}F(x)(1-F(x))$.
@@
Appendix C Calculation of ${\sf F}_{s}(\mathbf{z}_{k},\mathbf{h}_{k},\gamma)$ and ${\sf G}_{s}(\mathbf{z}_{k},\mathbf{h}_{k},\gamma)$
Given $\mathbf{z}_{k}$ in (34), the posterior distribution of $s_{k}$ given $\mathbf{z}_{k}$ is